OpenStudy (anonymous):

Solve by elimination method 15x-10y=5 15y=-6x+15x

OpenStudy (mertsj):

Rearrange the second equation so that it says -9x+15y=0 Then divide it by 3 to make the numbers smaller: -3x+5y=0 Now multiply it by 5 so that the x coefficients will be opposites. -15x +25y=0 Now add it to the top equation and solve.